# Singleton Park, Swansea SA2 8PP, UK

### v1

@@ 4.2 Additional Predictions
Aside from the dip at ${\tilde{\alpha}}^{\prime}t\approx-1$ discussed above, there are additional zeros of the amplitude at ${\tilde{\alpha}}^{\prime}u\approx-1$ (which is consistent with the position of the Veneziano dip) and ${\tilde{\alpha}}^{\prime}u\approx-0.4$. The latter, third zero is not present for all values of the energy. The positions of these additional minima are more dependent on the energy, as can be seen in figure 6. In particular, the right figure shows that the positions of these zeros are sensitive to the particular value of ${\tilde{\alpha}}^{\prime}s$. They are outside the range of the experimental data of Bromberg:1983he, and it would be an interesting feature to verify against other experiments.
@@
One can also see from the plots that the amplitude converges faster to the $s\to\infty$ limit when the angles are not small, i.e. when $|t|$ is also large, so we can already probe the high energy limit at ${\tilde{\alpha}}^{\prime}s\sim 10$.
@@ 5 Summary and Outlook
In this paper, we obtained predictions for the angular dependence of 4-point pion scattering amplitudes following the holographic approach of reference Armoni:2024gqv, in the specific case of the hard-wall AdS model. We compared our predictions with experimental data for $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering.  We focused primarily on the high-energy fixed-angle regime, though we also discussed the Regge regime.   Before summarising our findings, we briefly review some of the most important limitations of both our approach and the experimental data.
The ansatz for our pion scattering amplitudes, taken from reference Armoni:2024gqv, was constructed based on the work of Polchinski and Strassler Polchinski:2001tt.  Our predictions derived from this ansatz are restricted to tree-level scattering, with all calculations carried out at leading order in the supergravity approximation $\alpha^{\prime}/R_{5}^{2}\ll 1$.  With this approach, we successfully recovered the constituent counting rule and Regge behaviour. Since our predictions rely solely on the asympotically AdS region of the hard-wall geometry, which models QCD’s UV-behaviour via a CFT, they are expected to be reliable only in the high-energy fixed-angle limit.  In addition, note that our ansatz does not account for contributions from the WZW term (although none arise from vertices with two pions and an axial-vector meson, it would be worth checking whether vertices involving two pions and excited string states contribute), and quark masses are neglected.
On the other hand, the experimental data taken from reference Bromberg:1983he, against which we compared our predictions, is of limited reliability for pion scattering, as it was extracted from a baryon-pion scattering experiment.  Moreover, since additional phenomenological models were required to relate it to pion scattering, the validity of the extracted data beyond a qualitative level is limited.
@@
Our work can be extended and improved in several directions. Possible improvements include incorporating internal-space momentum contributions (e.g. the components $p_{w}^{(i)}$), a nontrivial dilaton profile, or quark masses. Further refinements will be especially valuable once more accurate pion-scattering data become available in the high-energy regime. Addressing some of these issues will likely require a better understanding of the Polchinski–Strassler proposal Polchinski:2001tt in string theory.
Our ansatz also shares similarities with that of reference Veneziano:2017cks, where a pion scattering amplitude based on infinitely many strings with varying tensions was proposed. It would be interesting to further develop this model and compare with ours and with the experimental data.
Recently, pion scattering amplitudes have also been computed using bootstrap techniques Guerrieri:2024jkn. It would be interesting to compare these results with our approach, if there is a common range of validity.
@@ Acknowledgements.
We are grateful to S. Sugimoto for insightful discussions and collaboration.  We thank M. Bianchi, Z. Komargodski, T. Sakai, M. Watanabe, M. Ward, N. Zenoni, R. Wendell, S. K. Sake, J. Barman, J. Nongmaithem, and X. U. Nguyen for discussions. DW was supported by an INFN postdoctoral fellowship.
@@ Appendix A Supplementary Figures
In Figure 9 are presented the fits to the 100 GeV data set of Bromberg:1983he. The fitting parameters are listed in Table 2 in the main text.
@@ Appendix B Evidence for 4-Point Contribution in Experimental Data
In this appendix, we plot some of the data for the component $I_{0}$ from the $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements reported in reference Armoni:2024gqv (for the definition of this component refer back to Section 3). For the same binned values of $s$, we superimpose the measurements from the 100 GeV and 175 GeV data sets. At a qualitative level, the behaviour at each binned value of $s$ appears to show strong resemblance. We take this to suggest that the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ includes a significant contribution from an underlying 4-point scattering mechanism. This observation was already previously noted in reference Bromberg:1983he; however, here we take the opportunity to plot the different data sets superimposed for direct comparison.
@@
Figure 10: Comparisons of the component $I_{0}$ from the $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements reported in reference Bromberg:1983he are shown for different binned $s$-values. The red data points correspond to the 100 GeV set, while the black data points correspond to the 175 GeV set. The data were extracted from Figure 29 in reference Bromberg:1983he. Note, however, that we were unable to recover all of the data points, and not all error bars could be obtained with sufficient precision; such instances were, however, rare.

### v2

@@ 4.2 Additional Predictions
Aside from the dip at ${\tilde{\alpha}}^{\prime}t\approx-1$ discussed above, there are additional zeros of the amplitude at ${\tilde{\alpha}}^{\prime}u\approx-1$ (which is consistent with the position of the Veneziano dip) and ${\tilde{\alpha}}^{\prime}u\approx-0.4$. The latter, third zero is not present for all values of the energy. The positions of these additional minima are more dependent on the energy, as can be seen in figure 6. In particular, the right figure shows that the positions of these zeros are sensitive to the particular value of ${\tilde{\alpha}}^{\prime}s$. They are outside the range of the experimental data of Bromberg and others (1984a), and it would be an interesting feature to verify against other experiments.
@@
One can also see from the plots that the amplitude converges faster to the $s\to\infty$ limit when the angles are not small, i.e. when $|t|$ is also large, so we can already probe the high-energy limit at ${\tilde{\alpha}}^{\prime}s\sim 10$.
@@ 5 Summary and Outlook
In this paper, we obtained predictions for the angular dependence of 4-point pion scattering amplitudes following the holographic approach of reference Armoni et al. (2025), in the specific case of the hard-wall AdS model. We compared our predictions with experimental data for $\pi^{+}\pi^{-}\to\pi^{+}\pi^{-}$ scattering.  We focused primarily on the high-energy fixed-angle regime, though we also discussed the Regge regime.   Before summarising our findings, we briefly review some of the most important limitations of both our approach and the experimental data.
The ansatz for our pion scattering amplitudes, taken from reference Armoni et al. (2025), was constructed based on the work of Polchinski and Strassler Polchinski and Strassler (2002).  Our predictions derived from this ansatz are restricted to tree-level scattering, with all calculations carried out at leading order in the supergravity approximation $\alpha^{\prime}/R_{5}^{2}\ll 1$.  With this approach, we successfully recovered the constituent counting rule and Regge behaviour. Since our predictions rely solely on the asympotically AdS region of the hard-wall geometry, which models QCD’s UV-behaviour via a CFT, they are expected to be reliable only in the high-energy fixed-angle limit.  In addition, note that our ansatz does not account for contributions from the WZW term (although none arise from vertices with two pions and an axial-vector meson, it would be worth checking whether vertices involving two pions and excited string states contribute), and quark masses are neglected.
On the other hand, the experimental data taken from reference Bromberg and others (1984a), against which we compared our predictions, is of limited reliability for pion scattering, as it was extracted from a baryon-pion scattering experiment.  Moreover, since additional phenomenological models were required to relate it to pion scattering, the validity of the extracted data beyond a qualitative level is limited.
@@
Our work can be extended and improved in several directions. Possible improvements include incorporating internal-space momentum contributions (i.e. the components $p_{w}^{(i)}$), a nontrivial dilaton profile, or quark masses. Further refinements will be especially valuable once more accurate pion-scattering data become available in the high-energy regime. Addressing some of these issues will likely require a better understanding of the Polchinski–Strassler proposal Polchinski and Strassler (2002) in string theory.
Our ansatz also shares similarities with that of reference Veneziano et al. (2017), where a pion scattering amplitude based on infinitely many strings with varying tensions was proposed. It would be interesting to further develop this model and compare with ours and with the experimental data.
Recently, pion scattering amplitudes have also been computed using bootstrap techniques Guerrieri et al. (2024). It would be interesting to compare these results with our approach, if there is a common range of validity.
@@ Acknowledgements.
We are grateful to S. Sugimoto for insightful discussions and collaboration.  We thank B. Pyszkowski for collaboration and discussion. We thank M. Bianchi, Z. Komargodski, T. Sakai, M. Watanabe, M. Ward, N. Zenoni, R. Wendell, S. K. Sake, J. Barman, J. Nongmaithem, and X. U. Nguyen for discussions. DW was supported by an INFN postdoctoral fellowship and the INFN project ST&FI “String Theory and Fundamental Interactions”.
@@ Appendix A Supplementary Figures
In Figure 9 are presented the fits to the 100 GeV data set of Bromberg and others (1984a). The fitting parameters are listed in Table 2 in the main text.
@@ Appendix B Evidence for 4-Point Contribution in Experimental Data
In this appendix, we plot some of the data for the component $I_{0}$ from the $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements reported in reference Bromberg and others (1984a) (for the definition of this component refer back to Section 3). For the same binned values of $s$, we superimpose the measurements from the 100 GeV and 175 GeV data sets. At a qualitative level, the behaviour at each binned value of $s$ appears to show strong resemblance. We take this to suggest that the process $\pi^{-}p\to\pi^{+}\pi^{-}n$ includes a significant contribution from an underlying 4-point scattering mechanism. This observation was already previously noted in reference Bromberg and others (1984a); however, here we take the opportunity to plot the different data sets superimposed for direct comparison.
@@
Figure 10: Comparisons of the component $I_{0}$ from the $\pi^{-}p\to\pi^{+}\pi^{-}n$ measurements reported in reference Bromberg and others (1984a) are shown for different binned $s$-values. The red data points correspond to the 100 GeV set, while the black data points correspond to the 175 GeV set. The data were extracted from Figure 29 in reference Bromberg and others (1984a). Note, however, that we were unable to recover all of the data points, and not all error bars could be obtained with sufficient precision; such instances were, however, rare.